A curve along the highway is to be designed for cars moving at 100 kilometers per hour. Find the angle of banking if the radius of the curve is 200 m.


1
Expert's answer
2021-05-20T11:50:16-0400

Let's apply the Newton's Second Law of Motion:


"Ncos\theta=mg, (1)""Nsin\theta=\dfrac{mv^2}{r}. (2)"

Dividing equation (2) by equation (1), we get:


"tan\theta=\dfrac{v^2}{gr},""\theta=tan^{-1}(\dfrac{v^2}{gr}),""\theta=tan^{-1}(\dfrac{(100\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s})^2}{9.8\ \dfrac{m}{s^2}\cdot200\ m})=21.5^{\circ}."